OpenStudy (anonymous):

36C3 evaluate the given expression

ganeshie8 (ganeshie8):

\(\large ^nC_r = \frac{n!}{(n-r)!r!}\)

ganeshie8 (ganeshie8):

\(\large ^{36}C_3 = \frac{36!}{(36-3)!3!}\)

ganeshie8 (ganeshie8):

\(\large ^{36}C_3 = \frac{36!}{(33)!3!}\)

ganeshie8 (ganeshie8):

\(\large ^{36}C_3 = \frac{36\times 35\times 34\times 33!}{(33)!3!}\)

ganeshie8 (ganeshie8):

\(33!\) cancels out on top and bottom

ganeshie8 (ganeshie8):

\(\large ^{36}C_3 = \frac{36\times 35\times 34}{3!}\)

ganeshie8 (ganeshie8):

\(\large ^{36}C_3 = \frac{36\times 35\times 34}{3\times 2}\)

ganeshie8 (ganeshie8):

simplify